OpenStudy (anonymous):

5/6x2/3 2/3tothesecondpower Please solve and show steps!

OpenStudy (anonymous):

I'm not sure if this is what you were meaning but: \[\frac{ 5 }{ 6 } \times \frac{ 2 }{ 3 } = \frac{ (5 \times 2) }{ (6 \times 3) } = \frac{ 10 }{ 18 } = \frac{ 5 }{ 9 }\] and for the second question \[(\frac{ 2 }{ 3 })^{2} = \frac{ 2 }{ 3 } \times \frac{ 2 }{ 3 } = \frac{ 2 \times 2 }{ 3 \times 3 } = \frac{ 4 }{ 9 }\]